OpenStudy (anonymous):

Please solve: start with 2 and multiply by -3 repeatedly

OpenStudy (kasim17):

What the hell dude? When are you support to stop?

OpenStudy (whpalmer4):

\[\begin{array}{cc} 1 & 2 \\ 2 & -6 \\ 3 & 18 \\ 4 & -54 \\ 5 & 162 \\ 6 & -486 \\ 7 & 1458 \\ 8 & -4374 \\ 9 & 13122 \\ 10 & -39366 \\ 11 & 118098 \\ 12 & -354294 \\ 13 & 1062882 \\ 14 & -3188646 \\ 15 & 9565938 \\ 16 & -28697814 \\ 17 & 86093442 \\ 18 & -258280326 \\ 19 & 774840978 \\ 20 & -2324522934 \\ 21 & 6973568802 \\ 22 & -20920706406 \\ 23 & 62762119218 \\ 24 & -188286357654 \\ 25 & 564859072962 \\ 26 & -1694577218886 \\ 27 & 5083731656658 \\ 28 & -15251194969974 \\ 29 & 45753584909922 \\ 30 & -137260754729766 \\ 31 & 411782264189298 \\ 32 & -1235346792567894 \\ 33 & 3706040377703682 \\ 34 & -11118121133111046 \\ 35 & 33354363399333138 \\ 36 & -100063090197999414 \\ 37 & 300189270593998242 \\ 38 & -900567811781994726 \\ 39 & 2701703435345984178 \\ 40 & -8105110306037952534 \\ 41 & 24315330918113857602 \\ 42 & -72945992754341572806 \\ 43 & 218837978263024718418 \\ 44 & -656513934789074155254 \\ 45 & 1969541804367222465762 \\ 46 & -5908625413101667397286 \\ 47 & 17725876239305002191858 \\ 48 & -53177628717915006575574 \\ 49 & 159532886153745019726722 \\ 50 & -478598658461235059180166 \\ \end{array}\]Need more? :-)

OpenStudy (kasim17):

Woah

OpenStudy (tommynaut):

If it's an equation you need, you could have it in the form \[y = 2\times(-3)^{x}\]

OpenStudy (whpalmer4):

Exercise for the reader: verify my answer (using pencil and paper) with @Tommynaut's formula for \(x = 50\) :-)

OpenStudy (whpalmer4):

Actually, we have a slight difference between our counting: I count "2" as the first number in the sequence, and he counts 2*(-3) = 6 as the first number in the sequence. Compare my 50 with what you get with \(x=49\) in his formula.

OpenStudy (kasim17):

-300

OpenStudy (tommynaut):

If the question really just wants you to do that and keep going on forever, maybe the best way to write it out would be \[\lim_{x \rightarrow \infty} y = 2(-3)^{x}\] And yes, I forget to mention that x=0 should be the start-point if 2 is included.

OpenStudy (tommynaut):

Oops, maybe the "y = " isn't necessary...

OpenStudy (kasim17):

Tommy, if this is for school it would make no sense if they want you to write it all out, and clearly he isn't even paying attention anymore.

OpenStudy (whpalmer4):

If you're trying to write the formula for a geometric series, in my experience that would be \[a_n = a*r^{n-1}\]where \(r\) is the common ratio, \(n\) is the number of the term you seek, and \(a = a_1\)

OpenStudy (tommynaut):

I don't think it is for school - I think he just made up some random question. If it was meant to be a geometric series question (which it sort of is) then I think he would have phrased it differently.

OpenStudy (whpalmer4):

If he (or any other reader who stumbles across this question) accidentally learns something by reading it, that isn't all bad, is it? :-)